TRUE or FALSE.


The sound intensity produced by a given sound source is directly proportional to the square of the distance from the source.

1
Expert's answer
2022-01-09T18:18:39-0500

We know that

Intensity of sound

"I=\frac{P}{A}"

"I=\frac{P}{4\pi r^2}"

I inversely proportional to squre of distance

"I\alpha\frac{1}{r^2}"

The sound intensity produced by a given sound source is directly proportional to the square of the distance from the source.

This statement is False

FALSE